``Before, solving the problem, I just want to give you the heads up. It's
Systems of Equation: Application. I can't find the specific title to relate
with my story problems. My chapter in the book is Chapter Eight, Solving Systems by Substiution, Additional. Hope it helps!
*Roberto decided to invest 20,000 in two accounts. One fo the accounts pays 5% interest and the other pays 8% interest. The interest he earned in one year was $1210. How much did he have invested in each account?*

Answer by Mathtut(3670)

let x and y be the amounts invested at 5 and 8% respectively
:
x+y=20000...............eq 1
.05x+.08y=1210..........eq 2
:
rewrite eq 1 to x=20000-y and plug this into eq 2
:
.05(20000-y)+.08y=1210
:
1000-.05y+.08y=1210
:
.03y=210
:
y=7000 dollars invested at 8%
:
x=20000-y=20000-7000=13000 invested at 5%